# Beyond Pairwise Learning-To-Rank At Airbnb

### v1

@@ 7. Stability Of Results
To enhance the stability of search results, we leverage the pairwise LTR logits generated by the first-stage ranker. Since this ranker is a univariate function that evaluates one listing at a time, its logits are inherently stable. For the second-stage ranker, we treat the all-pairwise LTR model as a residual adjustment layered on top of the first-stage pairwise LTR model (refer to Figure 5). By making the multivariate logit an incremental adjustment to the stable univariate logit, we effectively stabilize the overall search results. The experimental results section highlights the improvements in stability achieved using this approach.
@@ 8. Experimental Results
  • true-pairwise-avg: The final score for a listing is the average of the $(N-1)$ scores from true pairwise comparisons.
  • true-pairwise-gbt: The $(N-1)$ true pairwise comparisons are aggregated using the generalized Bradley-Terry model (Eq 9).
@@
  • all-pairwise-apfn: This uses the architecture in Figure 5, where the all-pairwise logit network is fed with superiority-similarity features from APFN.
@@
We start with offline experiments, using NDCG as the evaluation metric. Booked listings are assigned a relevance of 1, while not-booked listings are assigned a relevance of 0. Given the success of scaling laws for neural language models ((Kaplan et al., 2020),  (Hoffmann et al., 2022)), a frequently asked question is whether such scaling is applicable to our ranking models as well. Instead of investigating different LTR frameworks, can ranking performance be improved by simply scaling the model parameters of the baseline pairwise LTR model? In our first offline experiment, we study the effect of scaling the number of trainable model parameters (Figure  6).
@@
The results show that performance improvements obtained by LTR frameworks that better represent searcher behavior, cannot be achieved by simple scaling laws. All-pairwise LTR outperforms true-pairwise LTR, which in turn performs better than pairwise LTR, confirming the increase in accuracy predicted by theory. This creates a clear hierarchy: the more effectively a model captures interactions between listings, the better its ranking performance. Additionally, scaling beyond $2$ to $4$ million parameters causes overfitting across all variants, showing that indiscriminate scaling can harm performance.
Our second offline experiment focuses on scalability, where we examine the trade-off between cost and benefit for all-pairwise LTR. Figure  7 plots the relationship between the number of listings reranked by all-pairwise LTR in the second-stage ranker, and the resulting change in NDCG and latency. The growth in NDCG is sub-linear and flattens beyond a certain point, but the growth in latency continues to rise dramatically. Consequently, there is an optimal trade-off between benefit vs cost, and we settled at reranking $60$ listings in the second stage ranker.
@@
While Figure  6 clearly shows the superior performance of all-pairwise and true-pairwise LTR over pairwise LTR, in our fourth experiment we study a flip side—the uncertainty in model performance. Of the two varieties of uncertainty (Kendall and Gal, 2017)—aleatoric and epistemic—we focus on measuring epistemic uncertainty, since aleatoric uncertainty is independent of the LTR frameworks. Table 1 shows the range of NDCG obtained for the different LTR frameworks by varying the random seed that controls training data shuffling and neural network parameter initialization. The standard deviation in NDCB is around $2x$ to $6x$ higher compared to pairwise LTR. One possible reason for the higher epistemic uncertainty is that multivariate and bivariate models, with more inputs, can fit a broader range of parameters. However, averaging pairwise comparisons, as implemented in true-pairwise-avg, helps lower epistemic uncertainty.
@@
Figure 9. X-axis: The treatment models. Y-axis: % gain in uncancelled bookings.
@@
Our next experiment focuses on the multi-objective optimization discussed in Section 6. We pick all-pairwise-apn, the most effective ranking model identified so far, and train it with the weighted loss from Eq 11. We compare it with Airbnb’s production system, which used separate models to optimize bookings and trip quality. Results of the online AB test are highlighted in Table 2, which demonstrate the effectiveness of training all-pairwise LTR with the loss weighted by multi-objective.
@@
Table 2. Comparison of all-pairwise-apn trained with loss of Eq 11 vs active production ranker.
@@
Table 3. Comparing LTR frameworks under the SAT theorem. True-pairwise** is modified by Eq 8, followed by generalized Bradley-Terry model.
@@ 9. Conclusion
”You can have two Big Things, but not three.”
@@
Of course, the internet is onto something here. But we found a working compromise—combine two things from two places. In our research, all-pairwise LTR as the second-stage ranker, combined with pairwise LTR in the first stage, emerged as the clear winner. Table 3 provides a summary. The all-pairwise LTR model from Section 5, which uses superiority-similarity features and trained with the multi-objective approach from Section 6, was successfully deployed to $100\%$ of Airbnb searchers in early 2025. Prior to the rollout, we conducted an in-depth review of key business metrics, system performance, and its impact on developer productivity. Two questions were identified that remain unanswered: (1) Why all NDCG gains do not generalize equally to booking gains online, and what makes some flavors of NDCG gain more impactful than others? (2) Are there techniques that allow all-pairwise LTR to scale further to handle a larger number of listings? Our current research is focused on these questions.

### v2

@@ 7. Stability Of Results
To enhance the stability of search results, we leverage the pairwise LTR logits generated by the first-stage ranker. Since this ranker is a univariate function that evaluates one listing at a time, its logits are inherently stable. For the second-stage ranker, we treat the all-pairwise LTR model as a residual adjustment layered on top of the first-stage pairwise LTR model (refer to Figure 5). Adding the multivariate logit as a residual to the stable univariate logit significantly reduces the instability of search results. The experimental results section highlights the improvements in stability achieved using this approach.
@@ 8. Experimental Results
  • true-pairwise-avg: The final score for a listing is the average of the $N-1$ scores from true pairwise comparisons.
  • true-pairwise-gbt: The $N-1$ true pairwise comparisons are aggregated using the generalized Bradley-Terry model (Equation 9).
@@
  • all-pairwise-apfn: This uses the architecture in Figure 5, where the all-pairwise logit network is fed with superiority-similarity features from the APFN.
@@
We start with offline experiments, using NDCG as the evaluation metric. Booked listings are assigned a relevance of 1, while not-booked listings are assigned a relevance of 0. Given the success of scaling laws for neural language models (Kaplan et al., 2020)  (Hoffmann et al., 2022), a frequently-asked question is whether such scaling is applicable to our ranking models as well. Instead of investigating different LTR frameworks, can ranking performance be improved by simply scaling the model parameters of the pairwise LTR model? In our first offline experiment, we study the effect of scaling the number of trainable model parameters (Figure  6).
@@
The results show that performance improvements from LTR frameworks that better represent searcher behavior cannot be achieved by simple scaling laws. All-pairwise LTR outperforms true-pairwise LTR, which in turn performs better than pairwise LTR, confirming the increase in accuracy predicted by theory. This creates a clear hierarchy: the more effectively a model captures interactions between listings, the better its ranking performance. Additionally, scaling beyond $2$ to $4$ million parameters causes overfitting across all variants, showing that indiscriminate scaling can harm performance.
Our second offline experiment focuses on scalability, where we examine the trade-off between cost and benefit for all-pairwise LTR. Figure  7 plots the relationship between the number of listings reranked by all-pairwise LTR in the second-stage ranker, and the resulting change in NDCG and latency. The growth in NDCG is sub-linear and flattens beyond a certain point, but the growth in latency continues to rise dramatically. Consequently, there is an optimal trade-off between benefit vs. cost, and we settled at reranking $60$ listings in the second-stage ranker.
@@
While Figure  6 clearly shows the superior performance of all-pairwise and true-pairwise LTR over pairwise LTR, in our fourth experiment we study a flip side—the uncertainty in model performance. Of the two varieties of uncertainty (Kendall and Gal, 2017)—aleatoric and epistemic—we focus on measuring epistemic uncertainty, since aleatoric uncertainty is independent of the LTR frameworks. Table 1 shows the range of NDCG per LTR framework based on changes in random seed for data shuffling and network initialization. The standard deviation in NDCG is around 2x to 6x higher compared to pairwise LTR. One possible reason for the higher epistemic uncertainty is that multivariate and bivariate models, with more inputs, can fit a broader range of parameters. However, averaging pairwise comparisons, as implemented in true-pairwise-avg, helps lower epistemic uncertainty.
@@
Figure 9. X-axis: Treatment models. Y-axis: % gain in uncancelled bookings.
@@
Our next experiment focuses on the multi-objective optimization discussed in Section 6. We pick all-pairwise-apn, the most effective ranking model identified so far, and train it with the weighted loss from Equation 11. We compare it with Airbnb’s production system, which used separate models to optimize bookings and trip quality. Results of the online AB test are highlighted in Table 2, which demonstrate the effectiveness of training all-pairwise LTR with the loss weighted by multi-objective.
@@
Table 2. Comparison of all-pairwise-apn trained with loss of Equation 11 vs. active production ranker.
@@
Table 3. Comparing LTR frameworks under the SAT theorem. True-pairwise** is modified by Equation 8, followed by generalized Bradley-Terry model.
@@ 9. Conclusion
“You can have two Big Things, but not three.”
@@
Of course, the internet is onto something here. But we found a working compromise: combine two things from two places. In our research, all-pairwise LTR as the second-stage ranker, combined with pairwise LTR in the first stage, emerged as the clear winner. Table 3 provides a summary. The all-pairwise LTR model from Section 5, which uses superiority-similarity features and trained with the multi-objective optimization from Section 6, was successfully deployed to $100\%$ of Airbnb searchers in early 2025. Prior to the rollout, we conducted an in-depth review of key business metrics, system performance, and its impact on developer productivity. Two questions were identified that remain unanswered: (1) Why don’t all NDCG improvements lead to more bookings online—and what makes some types of NDCG gains matter more than others? (2) Are there techniques that allow all-pairwise LTR to scale further to handle a larger number of listings? Our current research is focused on these questions.